Topic: Normal Distribution Distribution & Inverse Normal Distribution

  1. Given that  X~N (4,9), find

a.) P(X > 8)

b.) P(4< X < 6)

c.) P(X≥3)



1
Expert's answer
2021-07-19T05:52:20-0400

Given "N(4, 3^2):\mu=4, \sigma=3"


a,


"P(X>9)=1-P(X\leq 9)="

"\approx1-P(Z\leq \dfrac{9-4}{3})\approx1-P(Z\leq 1.666667)"

"\approx0.047790"

b,


"P(4<X<6)=P(X<6)-P(X\leq 4)="

"\approx P(Z< \dfrac{6-4}{3})-P(Z\leq \dfrac{4-4}{3})"

"\approx P(Z\leq 0.666667)-P(Z\leq 0)"

"\approx0.747507-0.5\approx0.247507"

c,


"P(X\geq3)=1-P(X<3)="

"\approx1-P(Z< \dfrac{3-4}{3})\approx1-P(Z\leq -0.333333)"

"\approx0.630559"